OpenStudy (anonymous):

Find P(rolling 2 or 5) with one number cube. A.1/2 B.1/3 C.1/6 D.1

OpenStudy (anonymous):

The answer is 1/3.

OpenStudy (anonymous):

2 options/6 possibilities = 2/6 = 1/3

OpenStudy (anonymous):

options is probably not a good term, but you get the idea.

OpenStudy (anonymous):

Yeah , i know what you mean .

OpenStudy (anonymous):

answer: 1/3 because 2 options divided 6 possibilities would equal 2/6 = 1/3 @Nanoman is right

OpenStudy (anonymous):

= P (2) + P (5) = 1/6 + 1/6 = 2/6 = 1/3